OpenStudy (anonymous):

help Which of the following points lies on the circle whose center is at the origin and whose radius is 13? (-5, 12) (√13,√13) (6, -7)

OpenStudy (anonymous):

the equation of the circle is \[x^2+y^2=13\]

OpenStudy (anonymous):

now check which one is satisfying the equation

OpenStudy (anonymous):

so its the 2nd one

OpenStudy (anonymous):

sorry \[x^2+y^2=13^2\]

OpenStudy (anonymous):

now say which is satisfying it..

OpenStudy (anonymous):

so the 3rd one

OpenStudy (anonymous):

1st one

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

\[(-5)^2+12^2=13^2\]

OpenStudy (anonymous):

welcome :)